# False Discovery Rate and Localizing Power

### v1

@@ 2.8 Direction and localization errors
As noted above, fdr does not provide strong control over the error rate, which means that the amount of errors in any subset of tests is not guaranteed to be within the desired level. For two-tailed tests, this becomes problematic if the researcher wishes to draw conclusions pertaining to the direction of effects — that is, inferring about each tail separately. Likewise, this becomes problematic if the researcher wishes to draw conclusions about the amount of errors within specific regions after fdr has been applied globally. To illustrate that, consider these two examples:
@@
An investigation into the association between cortical thickness (measured at every vertex of a surface representation of the brain) and a behavioral measure involves a sample of hundreds of individuals. Acknowledging that effects could be manifest in either direction, and pursuing statistical rigor, the researcher employs two-tailed p-values, with adjustments for multiple testing using an fdr. The map of test statistics is then masked at the significant level $q=0.01$, thus retaining for visualization only the significant vertices. The analysis uncovers large areas in the frontal and parietal lobes with a positive relationship to the behavioral measure, as hypothesized. However, it also reveals a negative association spanning broad portions of the fusiform gyrus, as well as vertices showing positive or negative associations scattered across the cortex. The researcher conducting the analysis writes a report detailing the novel negative findings, emphasizing the stringent methodologies that incorporated two-tailed tests and a strict fdr significance threshold, well below the conventional level $q=0.05$, and concludes (erroneously) that only about 1% of the negative associations are expected to be false positives.
@@
In a study of a compound for reducing the severity of relapses in multiple sclerosis (ms), white matter hyperintensities (wmh) are analyzed voxelwise. Subjects are randomly allocated into active compound and placebo groups. Based on prior knowledge and on random population sampling, normality is assumed and a two-sided group comparison is conducted using a two-sample $t$-test. This produces two-tailed p-values, calculated from the Student’s $t$-distribution. Control for multiple testing across voxels is achieved using fdr at $q=0.05$. After determining the fdr threshold, the inverse cumulative distribution function of the Student’s $t$-distribution is used to derive the critical threshold $t_{\textsc{fdr}}$. Given the two-tailed nature of the test, symmetrical thresholds $+t_{\textsc{fdr}}$ and $-t_{\textsc{fdr}}$ are set in an image visualization tool. The results show a notable reduction in the periventricular wmh load in the group that used the active compound, as expected (and hoped), as well as mixed results in various juxtacortical regions, albeit with mostly increases in wmh load in these areas. In the report, the investigator concludes (erroneously) that the drug may amplify the wmh load in juxtacortical areas given that, as fdr was used, no more than 5% of the observed changes are false positives.
@@
At first, it may seem that the researchers in both examples did everything correctly, going as far as using two-tailed tests and ensuring that the multiplicity of tests would be addressed rigorously. A casual reader of the respective study reports may find no issue. Yet, where it comes to statements about the direction or localization of effects, the interpretation is not that, among the findings in the negative direction (Example 1) or in a particular region (Example 2), the proportion of false positives is within $q$, but only that the proportion of false positives among all tests that were subjected to fdr is within $q$. The actual, realized (but unknown) fdr is not guaranteed to be within 1% for the negative (or positive) direction in Example 1, nor or 5% within the juxtacortical white matter (or any other region) in Example 2, and can in both cases be anywhere between 0 and 100% in the reported direction or region.
@@
These issues are not inherent to fdr, although they are a consequence of how fdr operates on data and its properties; the cause of the problem is the attempt to make inferences on subsets of the data over which fdr was applied. Localization errors can be avoided by refraining from making statements or from taking actions that could refer to the amount of errors in specific regions when fdr is controlled only globally, or by doing separate fdr analyses within each subset (Efron, 2008). The same holds to avoid direction errors. Direction errors can also be avoided by partitioning the statistical map into two disjoint sets: one with the positive test statistics, another with the negative test statistics, then proceeding to fdr correction in each subset using their two-tailed p-values. This is expected to work because under the null hypothesis, the partitioning using the sign of the test statistic is independent of the two-tailed p-values; the latter are distributed uniformly between 0 and 1 and, in either side of the map, more extreme test statistics (stronger effects) imply lower p-values. Thus, the requirements for fdr, using either bh or bky, are satisfied. The fdr can then be controlled at the desired level $q$ for the positive and for the negative sides of the map. We verify this novel proposition in the next section, along with other methods.
@@ 3.1 Synthetic data
To assess the realized (empirical) fdr across procedures we used synthetic data formed by 2000 tests, which could represent voxels in a group-level result image. For each, a test statistic $z$ was simulated following a normal distribution with zero mean and unit variance. Five different scenarios were considered for independent test statistics: (i) complete null, with no synthetic effects added; (ii) positive effects only, in which 25% of the tests had their $z$ statistic increased by 3; (iii) negative effects only, in which 25% of the tests had their $z$ statistic decreased by 3; (iv) both positive and negative effects, balanced, in which 25% of the tests had their $z$ statistic increased by 3, another 25% had their $z$ statistic decreased by 3; and (v) both positive and negative effects, unbalanced, in which 10% of the tests had their $z$ statistic increased by 3, another 40% had their $z$ statistic decreased by 3. Another five scenarios (vix) were considered; these were identical to the first five except that instead of independent tests, the 2000 test statistics were made non-independent, with a correlation of 0.25. In each scenario, p-values were computed by referring to the normal cumulative distribution function. A summary of the scenarios is shown in Table 1.
Table 1: Scenarios used to evaluate the different fdr approaches with synthetic data. In each scenario, 2000 test statistics were simulated and subjected to fdr procedures. The process was repeated 2000 times.
@@
Simulations for these 10 scenarios were repeated 2000 times. Each was assessed with both bh and bky. For this simulated data evaluation, we use $q=0.05$. Four different strategies were considered:
@@
  • Canonical, in which adjustment is applied to one-tailed p-values in one direction, and to one-tailed p-values in the opposite direction, that is, applied separately to $p$ and to $1-p$. This procedure produces two maps of adjusted p-values, one for each direction; alternatively, it produces two thresholds, one for each direction.
@@
  • Combined, in which adjustment is applied simultaneously to one-tailed p-values in the two directions, i.e., on the set formed by $p$ and $1-p$. This procedure produces two maps of adjusted p-values, one for each direction; alternatively, it produces one threshold, to be applied to both directions.
@@
  • Two-tailed, in which two-tailed p-values are subjected to adjustment. This procedure produces one map of adjusted p-values; alternatively, it produces one threshold.
@@
  • Split-tails, in which the tests are separated into two disjoint subsets, one with positive ($z>0$) and another with negative ($z<0$) statistics; then the two-tailed p-values within each subset are subjected to adjustment. For the cases in which the test statistic is zero ($z=0$), the uncorrected two-tailed p-value is 1, and so is the adjusted p-value. This procedure produces one map of adjusted p-values; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics.
For each of the ten scenarios, and for each of the four adjustment methods, which were run with each of bh and bky (80 total sets), the empirical (observed) fdr was calculated. Confidence intervals (95%) were calculated with the Wald method.
@@ 3.2 Real data
To assess these four different strategies for two-tailed fdr with real data, as well as to provide an empirical comparison between bh and bky, we use a statistical map originally produced as part of the Neuroimaging Analysis Replication and Prediction Study (narpsBotvinik-Nezer et al., 2020). The narps was an effort to assess the variability of results among different analysis teams analyzing the same imaging dataset. In the study, the teams were provided with the same raw functional magnetic resonance imaging (fmri) data, collected during a gambling task. Each team analyzed the data to investigate predefined hypotheses about brain activation related to gains and losses, and employed diverse analysis workflows, with no two teams choosing identical pipelines. We use (for no particular reason) the maps produced by team 9U7M for narps hypothesis 1. Note that the specifics of the hypothesis are not relevant for the present analysis; consult Botvinik-Nezer et al. (2020) for details about data acquisition, the predefined hypotheses, and the overall narps methodology and its data provenance. For this real data evaluation, we use $q=0.05$.
@@ 4.1 Synthetic data
Table 2 summarizes the results for the analyses using synthetic data across the 10 different scenarios that involved varying proportions of true positive and negative effects, as well as dependencies among tests. For a given test, the ideal result is that the confidence interval includes the chosen $q=0.05$ value (represented in percentage, i.e., 5%).
Table 2: Comparison of the false discovery proportions (%) among different fdr methods and correction strategies, when false positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The expected proportion of false discoveries is $q\cdot 100\%=5\%$. Confidence intervals (95%) are shown between parentheses.
@@
Scenario
Canonical
Combined
Two-tailed
Split-tails
Canonical
Combined
Two-tailed
Split-tails

### v2

@@ 2.6 Localizing power
The ability to localize effects within parts of an image requires that the error rate is controlled within that part. Not having localizing power is problematic, for example, if the researcher wishes to draw conclusions within specific regions after fdr has been applied globally. Likewise, this becomes problematic if the researcher wishes to draw conclusions pertaining to the direction of effects — that is, inferring about each tail separately — after applying fdr on two-tailed p-values. To illustrate that, consider these two hypothetical examples:
@@
An investigation into the association between cortical thickness (measured at every vertex of a surface representation of the brain) and a behavioral measure involves a sample of hundreds of individuals. Acknowledging that effects could be manifest in either direction, and pursuing statistical rigor, the researcher employs two-tailed p-values, with adjustments for multiple testing using an fdr. The map of test statistics is then masked at the significant level $q=0.01$, thus retaining for visualization only the significant vertices. The analysis uncovers large areas in the frontal and parietal lobes with a positive relationship to the behavioral measure, as hypothesized. However, it also reveals a negative association spanning broad portions of the fusiform gyrus, as well as vertices showing positive or negative associations scattered across the cortex. The researcher conducting the analysis writes a report detailing the novel findings in the negative direction, emphasizing the stringent methodologies that incorporated two-tailed tests and a strict fdr significance threshold, well below the conventional level $q=0.05$, and concludes (erroneously) that only about 1% of the negative associations are expected to be false positives.
@@
In a study of a compound for reducing the severity of relapses in multiple sclerosis (ms), white matter hyperintensities (wmh) are analyzed voxelwise. Subjects are randomly allocated into active compound and placebo groups. Based on prior knowledge and on random population sampling, normality is assumed and a two-sided group comparison is conducted using a two-sample $t$-test. This produces two-tailed p-values calculated from the Student’s $t$-distribution. Control for multiple testing across voxels is achieved using fdr at $q=0.05$. After determining the fdr threshold, the inverse cumulative distribution function of the Student’s $t$-distribution is used to derive the critical threshold $t_{\textsc{fdr}}$. Given the two-tailed nature of the test, symmetrical thresholds $t_{\textsc{fdr}}$ and $t_{\textsc{fdr}}$ are set in an image visualization tool. The results show a notable reduction in the periventricular wmh load in the group that used the active compound, as expected (and hoped), as well as mixed results in various juxtacortical regions, albeit with mostly increases in wmh load in these areas. In the report, the investigator concludes (erroneously) that the drug may amplify the wmh load in juxtacortical areas given that, as fdr was used, no more than 5% of the observed changes are false positives.
@@
At first, it may seem that the researchers in both examples did everything correctly, going as far as using two-tailed tests and ensuring that the multiplicity of tests would be addressed rigorously. A casual reader of the respective study reports may find no issue. Yet, where it comes to statements about the direction or localization of effects, the interpretation is not that, among the findings in the negative direction (Example 1) or in a particular region (Example 2), the proportion of false positives is within $q$, but only that the proportion of false positives among all tests that were subjected to fdr is within $q$. The actual, realized (but unknown) fdr is not guaranteed to be within 1% for the negative (or positive) direction in Example 1, nor 5% within the juxtacortical white matter (or any other region) in Example 2, and can in both cases be anywhere between 0 and 100% in the reported direction or region.
@@
These issues are a consequence of the lack of the subsetting property. Localization errors can be avoided by refraining from making statements or taking actions that could refer to the amount of errors in specific regions when fdr is controlled only globally, or by doing separate fdr analyses within each subset (Efron, 2008). The same holds to avoid direction errors. Direction errors can also be avoided, we assert, by partitioning the statistical map into two disjoint sets: one with the positive test statistics, another with the negative test statistics, then proceeding to fdr correction in each subset using their two-tailed p-values. This is expected to work because under the null hypothesis, the partitioning using the sign of the test statistic is independent of the two-tailed p-values; the latter are distributed uniformly between 0 and 1 and, in either side of the map, more extreme test statistics (stronger effects) imply lower p-values. Thus, the requirements for fdr, using either bh or bky, are satisfied. The fdr can then be controlled at the desired level $q$ for the positive and for the negative sides of the map. We verify this in the next section, along with other methods.
@@ 3.1 Synthetic data
To assess the realized (empirical) fdr across procedures we used synthetic data formed by 2000 tests, which could represent voxels in a group-level result image. For each, a test statistic $z$ was simulated following a normal distribution with zero mean and unit variance. Five different scenarios were considered for independent test statistics: (i) complete null, with no synthetic effects added; (ii) positive effects only, in which 25% of the tests had their $z$ statistic increased by 3; (iii) negative effects only, in which 25% of the tests had their $z$ statistic decreased by 3; (iv) both positive and negative effects, balanced, in which 25% of the tests had their $z$ statistic increased by 3, another 25% had their $z$ statistic decreased by 3; and (v) both positive and negative effects, unbalanced, in which 10% of the tests had their $z$ statistic increased by 3, another 40% had their $z$ statistic decreased by 3. Another five scenarios (vix) were considered; these were identical to the first five except that instead of independent tests, the 2000 test statistics were made non-independent, with a correlation of 0.25. In each scenario, uncorrected p-values were computed by referring to the normal cumulative distribution function. A summary of the scenarios is shown in Table 1.
Table 1: Scenarios used to evaluate the different fdr approaches with synthetic data. In each scenario, 2000 test statistics were simulated and subjected to fdr procedures. The process was repeated 2000 times. Scenarios iv contain independent tests; tests in scenarios vix have nonzero dependency, as indicated by the “Correlation” column.
@@
Simulations for these 10 scenarios were repeated (realized) 2000 times. Each was assessed with both bh and bky. For this simulated data evaluation, we use $q=0.05$. Four different strategies were considered, described below and summarized in Table 2. Hence, we have 80 sets (10 scenarios, 2 fdr procedures, and 4 strategies) of results, each based on 2000 realizations.
Table 2: Summary of the four strategies considered for running fdr.
Strategy Inputs to fdr Calls to fdr
Canonical $\{p^{\text{one}}_{i}\}$, $\{1-p^{\text{one}}_{i}\}$ Two, each with $V$ tests
Combined $\{p^{\text{one}}_{i},1-p^{\text{one}}_{i}\}$ One, with $2V$ tests
Two-tailed $\{p^{\text{two}}_{i}\}$ One, with $V$ tests
Split-tails $\{p^{\text{two}}_{i}:z_{i}>0\}$, $\{p^{\text{two}}_{i}:z_{i}<0\}$ Two, with a total of $V$ tests
For each synthetic voxel $i\in\{1,2,\ldots,V\}$, $p^{\text{one}}_{i}$ refers to their one-tailed p-values, $p^{\text{two}}_{i}$ refers to their two-tailed p-values; $z_{i}$ are their corresponding test statistics (for example, it could be the result of a signed statistical test, such as the $t$-test).
@@
  • Canonical, in which adjustment is applied to one-tailed p-values in one direction, and to one-tailed p-values in the opposite direction, that is, applied separately to $p$ and to 1p1-p. This procedure produces two complete maps of adjusted p-values, one for each direction; alternatively, it produces two thresholds, one for each direction.
@@
  • Combined, in which adjustment is applied simultaneously to one-tailed p-values in the two directions, i.e., on the set formed by $p$ and 1p1-p. This procedure produces two complete maps of adjusted p-values, one for each direction; alternatively, it produces one threshold, to be applied to both directions.
@@
  • Two-tailed, in which two-tailed p-values are subjected to adjustment. This procedure produces one complete map of adjusted p-values; alternatively, it produces one threshold.
@@
  • Split-tails, in which the tests are separated into two disjoint subsets, one with positive ($z>0$) and another with negative ($z<0$) statistics; then the two-tailed p-values within each subset are subjected to adjustment. For the cases in which the test statistic is zero ($z=0$), the uncorrected two-tailed p-value is 1, and so is the adjusted p-value; such pin risk is extremely unlikely with continuous test statistics. After the two disjoint sets are joined, this procedure produces one complete map of adjusted p-values; alternatively, it produces two thresholds, one for positive test statistics, one for negative test statistics.
Counting true and false positives
For each realization, true and false positives are considered in relation to the true signal; for directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; these definitions are formalized in Table 3. The proportion of false discoveries within a realization is the number of false discoveries divided by the total number of discoveries.
Assessing power
The power of each of the fdr methods and the four different strategies is the ratio between the number of true discoveries by the number of tests in which true signal was added, computed for reach realization. For directional inferences (positive or negative), these refer to the results in the positive and negative sides of the statistical map; for non-directional (both sides), the directions of the test statistic and of the true signal are ignored; as described in Table 3. Thus, power here represents a distinct quantity in relation to the usual statistical power. Here it represents the sensitivity of a multiple testing procedure judged by its ability to recover a number of tests that contain true signal, that is, tests in which the null hypothesis is false.
Table 3: Counting of tests that meet criteria for containing a true synthetic effect in a given direction, for having the sign of the test statistic matching the direction of the signal, and being statistically significant after an fdr-controlling procedure. These counts are used to calculate the empirical proportion of false discoveries and empirical power with synthetic data.
Direction Counts
Both sides:
[…]
True discoveries $\{i:(p^{\text{adj}}_{i}\leqslant q)\text{ and }(\top^{+}_{i}\text{ or }\top^{-}_{i})\}$
False discoveries $\{i:(p^{\text{adj}}_{i}\leqslant q)\text{ and }\text{not }(\top^{+}_{i}\text{ or }\top^{-}_{i})\}$
Positive side:
[…]
Negative side:
True effects $\{i:\top^{-}_{i}\}|$
[…]
For each synthetic voxel $\in\{1,2,\ldots,V\}, i+\top^{+}_{i}$ and $\top^{-}_{i}$ indicate whether a true positive or true negative effect was added to that voxel, $z_{i}$ represents the corresponding test statistic at that voxel, and $p^{\text{adj}}_{i}\leqslant q$ represents a significant adjusted p-value at that voxel after an fdr procedure. For each realization, the counts are the sizes (cardinalities) of each of the sets. The empirical fdr within a realization is the number of false discoveries observed in a given direction divided by the total number of discoveries in that direction, whereas power is the number of true discoveries in a given direction divided by the number of tests with true effects in that direction. For computations that refer to both sides, the direction is ignored. Averaging the proportion of false discoveries or the calculated power across realizations yields the results shown in Tables 4 and 5, respectively.
The overall empirical fdr (observed false discovery proportion) and power were computed by averaging across realizations. Confidence intervals (95%) were calculated with the Wald method (Brown et al., 2001).
@@ 3.2 Real data
To assess these four different strategies for two-tailed fdr with real data, as well as to provide an empirical comparison between bh and bky, we use a statistical map originally produced as part of the Neuroimaging Analysis Replication and Prediction Study (narpsBotvinik-Nezer et al., 2020). The narps was an effort to assess the variability of results among different analysis teams analyzing the same imaging dataset. In the study, the teams were provided with the same raw functional magnetic resonance imaging (fmri) data, collected during a gambling task. Each team analyzed the data to investigate predefined hypotheses about brain activation related to gains and losses, and employed diverse analysis workflows, with no two teams choosing identical pipelines. We use (for no particular reason) the maps produced by team 9U7M for narps Hypothesis 1. Note that the specifics of the hypothesis are not relevant to the present analysis; consult Botvinik-Nezer et al. (2020) for details about data acquisition, the predefined hypotheses, the overall narps methodology, and its data provenance. For this real data evaluation, we use $q=0.05$.
@@ 4.1 Synthetic data
Tables 4 and 5 summarize the results for the analyses using synthetic data across the 10 different scenarios that involved varying proportions of true positive and negative effects, as well as dependencies among tests. For a given test, the ideal result is that the confidence interval includes the chosen $q=0.05$ value (represented in the Table in percentage, i.e., 5%).
Table 4: (page 4.1) Comparison of the false discovery proportions (%) among different fdr methods and correction strategies, when false positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
Table 5: (page 4.1) Comparison of power (%) among different fdr methods and correction strategies, when true positives are counted considering one side (positive or negative) or both sides of the statistical map. These results are based on synthetic data. The nominal proportion of false discoveries is $q\cdot 100\%=5\%$; their corresponding confidence intervals (95%) are shown between parentheses.
@@
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails
Both sides:
i 9.5 (8.2–10.8) 5.1 (4.1–6.0) 5.1 (4.1–6.0) 9.6 (8.3–10.8) 9.5 (8.2–10.8) 4.8 (3.8–5.7) 4.8 (3.8–5.7) 9.6 (8.3–10.8)
$i$ 3.8 (3.7–3.8) 3.7 (3.7–3.8) 3.7 (3.7–3.8) 3.0 (3.0–3.1) 4.7 (4.6–4.7) 4.1 (4.1–4.2) 4.5 (4.5–4.6) 4.3 (4.3–4.4)
iii 4.0 (3.9–4.0) 4.0 (3.9–4.0) 4.0 (3.9–4.0) 3.2 (3.2–3.3) 4.9 (4.9–4.9) 4.2 (4.2–4.3) 4.7 (4.6–4.7) 4.6 (4.5–4.6)
iv 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 3.2 (3.2–3.2) 3.2 (3.2–3.2) 4.3 (4.3–4.3) 4.3 (4.3–4.3)
v 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.6 (2.6–2.6) 2.3 (2.3–2.3) 3.7 (3.7–3.7) 3.1 (3.1–3.2) 4.3 (4.2–4.3) 4.2 (4.2–4.3)
vi 7.7 (6.5–8.9) 4.5 (3.5–5.4) 4.5 (3.5–5.4) 5.4 (4.4–6.4) 8.1 (6.9–9.2) 4.0 (3.2–4.9) 4.0 (3.2–4.9) 4.7 (3.8–5.6)
vii 3.9 (3.6–4.3) 3.8 (3.5–4.1) 3.8 (3.5–4.1) 2.5 (2.3–2.7) 5.3 (4.9–5.8) 4.2 (3.9–4.6) 4.6 (4.3–5.0) 3.7 (3.3–4.0)
viii 4.4 (4.1–4.8) 4.2 (3.9–4.5) 4.2 (3.9–4.5) 2.9 (2.6–3.1) 6.0 (5.6–6.5) 4.6 (4.3–4.9) 5.0 (4.7–5.3) 4.0 (3.7–4.3)
ix 3.7 (3.5–3.9) 2.7 (2.5–2.8) 2.7 (2.5–2.8) 2.9 (2.7–3.0) 4.8 (4.5–5.1) 3.2 (3.1–3.4) 4.3 (4.1–4.5) 4.6 (4.3–4.8)
x 3.3 (3.1–3.4) 2.5 (2.4–2.7) 2.5 (2.4–2.7) 2.4 (2.3–2.6) 4.9 (4.6–5.2) 3.2 (3.0–3.3) 4.3 (4.1–4.4) 4.3 (4.1–4.5)
Positive side:
i 5.3 (4.4–6.3) 2.9 (2.2–3.7) 2.9 (2.2–3.7) 5.4 (4.4–6.4) 5.3 (4.4–6.3) 2.9 (2.1–3.6) 2.9 (2.1–3.6) 5.4 (4.4–6.4)
$i$ 3.8 (3.7–3.8) 1.9 (1.9–1.9) 1.9 (1.9–1.9) 3.0 (3.0–3.0) 4.7 (4.6–4.7) 2.1 (2.1–2.1) 2.3 (2.3–2.4) 4.3 (4.3–4.4)
iii 4.8 (3.8–5.7) 99.7 (99.4–99.9) 99.7 (99.4–99.9) 5.9 (4.9–7.0) 3.5 (2.7–4.3) 99.9 (99.8–100.0) 100.0 (99.9–100.0) 5.0 (4.0–5.9)
iv 2.5 (2.5–2.6) 2.5 (2.5–2.5) 2.5 (2.5–2.5) 2.5 (2.5–2.6) 3.1 (3.1–3.1) 3.1 (3.1–3.1) 4.2 (4.2–4.2) 4.2 (4.2–4.2)
v 2.5 (2.5–2.6) 6.0 (5.9–6.1) 6.0 (5.9–6.1) 3.6 (3.5–3.7) 2.7 (2.6–2.7) 7.3 (7.2–7.4) 9.8 (9.7–9.9) 4.4 (4.3–4.5)
vi 4.3 (3.5–5.2) 2.6 (1.9–3.4) 2.6 (1.9–3.4) 3.0 (2.3–3.8) 3.9 (3.1–4.7) 2.1 (1.5–2.8) 2.1 (1.5–2.8) 2.4 (1.7–3.0)
vii 3.6 (3.3–3.9) 1.8 (1.7–2.0) 1.8 (1.7–2.0) 2.3 (2.2–2.5) 4.7 (4.3–5.0) 1.9 (1.7–2.1) 2.2 (2.0–2.4) 3.2 (3.0–3.4)
viii 3.6 (2.8–4.4) 70.1 (68.1–72.1) 70.1 (68.1–72.1) 3.1 (2.4–3.9) 3.8 (2.9–4.6) 70.5 (68.5–72.5) 74.6 (72.7–76.5) 2.8 (2.1–3.5)
ix 2.4 (2.2–2.5) 1.9 (1.8–2.0) 1.9 (1.8–2.0) 2.0 (1.9–2.1) 3.1 (2.9–3.4) 2.4 (2.3–2.6) 3.4 (3.2–3.5) 3.4 (3.2–3.6)
x 2.2 (2.0–2.4) 3.6 (3.4–3.8) 3.6 (3.4–3.8) 2.4 (2.3–2.6) 2.7 (2.4–2.9) 4.7 (4.4–4.9) 6.3 (6.0–6.6) 3.3 (3.1–3.5)
Negative side:
i 4.5 (3.6–5.5) 2.5 (1.8–3.2) 2.5 (1.8–3.2) 4.5 (3.6–5.5) 4.3 (3.5–5.2) 2.1 (1.5–2.7) 2.1 (1.5–2.7) 4.3 (3.5–5.2)
$i$ 4.2 (3.3–5.1) 99.8 (99.5–100.0) 99.8 (99.5–100.0) 5.8 (4.8–6.8) 4.0 (3.1–4.9) 99.9 (99.8–100.0) 100.0 (99.9–100.0) 5.7 (4.7–6.7)
iii 4.0 (3.9–4.0) 2.1 (2.1–2.2) 2.1 (2.1–2.2) 3.2 (3.2–3.3) 4.9 (4.8–4.9) 2.3 (2.2–2.3) 2.5 (2.5–2.5) 4.6 (4.5–4.6)
iv 2.7 (2.7–2.7) 2.7 (2.6–2.7) 2.7 (2.6–2.7) 2.7 (2.6–2.7) 3.3 (3.2–3.3) 3.3 (3.2–3.3) 4.4 (4.3–4.4) 4.4 (4.3–4.4)
v 2.6 (2.6–2.6) 1.7 (1.7–1.7) 1.7 (1.7–1.7) 2.0 (2.0–2.0) 3.9 (3.9–3.9) 2.0 (2.0–2.1) 2.8 (2.7–2.8) 4.2 (4.2–4.2)
vi 3.4 (2.6–4.1) 1.8 (1.2–2.4) 1.8 (1.2–2.4) 2.4 (1.7–3.0) 4.2 (3.3–5.0) 1.9 (1.3–2.5) 1.9 (1.3–2.5) 2.4 (1.7–3.0)
vii 2.8 (2.1–3.5) 70.4 (68.4–72.4) 70.4 (68.4–72.4) 2.4 (1.7–3.0) 3.0 (2.3–3.8) 73.5 (71.6–75.4) 76.3 (74.4–78.2) 2.9 (2.2–3.6)
viii 3.7 (3.5–4.0) 2.0 (1.8–2.1) 2.0 (1.8–2.1) 2.5 (2.3–2.6) 5.3 (5.0–5.7) 2.3 (2.2–2.5) 2.7 (2.5–2.9) 3.7 (3.5–4.0)
ix 2.7 (2.5–2.9) 2.2 (2.0–2.3) 2.2 (2.0–2.3) 2.3 (2.1–2.4) 3.5 (3.3–3.8) 2.7 (2.5–2.8) 3.7 (3.4–3.9) 3.8 (3.5–4.0)
x 2.6 (2.4–2.7) 1.5 (1.4–1.7) 1.5 (1.4–1.7) 1.8 (1.6–1.9) 4.1 (3.9–4.4) 1.9 (1.8–2.0) 2.7 (2.5–2.9) 3.6 (3.4–3.8)
Benjamini–Hochberg Benjamini–Krieger–Yekutieli
Scenario Canonical Combined Two-tailed Split-tails Canonical Combined Two-tailed Split-tails